Why does 2,3-dichlorobutane have only 3 stereoisomers?
Answer #1

Theoretically a compound with 2 chiral centers will have 2^n = 2^2 = 4 stereoisomers. However, because the two chiral centers have identical substituents, one of the stereoisomers is meso and is superimposable on its mirror image, so there are only three stereoisomers: RR, SS, and RS(meso). Such that existence of a stereoisomer as meso, the number of possible isomers gets reduced to 1 less, that is we will have 3 stereoisomers in 1,2-dichlorobutane.